The length of a rectangle is 3 more than twice the width. The area is 44 square inches. Find the dimension
Answer by fractalier(6550)

A = LW = 44
and
L = 3 + 2W
now substitute this into the first equation...
(3+2W)W=44
3W+2W^2=44
Rearranging
2W^2 + 3W - 44 = 0
Factoring we get
(2W + 11)(W - 4) = 0
so that
W = 4 inches
L = 11 inches